OpenStudy (anonymous):

(-2)(-4)(-3) - (-3)(-4) = ? How do I solve and what is the solution?

OpenStudy (anonymous):

-36

OpenStudy (anonymous):

-2*-4=8 8*-3=-24 -3*-4=12 -24+-12=-36